OpenStudy (anonymous):

use substitution to determine whether the point is on line y=3x+7; A(2,13)

jimthompson5910 (jim_thompson5910):

We have the point (2,13) That means (x,y) = (2,13) which breaks down to x = 2 and y = 13

jimthompson5910 (jim_thompson5910):

To see if (2,13) is on the line, we just plug in (x,y) = (2,13) If we get a true equation, then the point is indeed on the line. If we get a false equation, then the point does not lie on the line (no matter how close it looks) y = 3x+7 13 = 3*2+7 ... plug in x = 2 and y = 13 13 = 6+7 13 = 13 ... true equation So this proves that (2,13) is on the line y=3x+7

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

you're welcome